Taylor United is considering overhauling its equipment to meet increased demand for its product. The cost of equipment overhaul is $4.15 million, plus $222,493.00 in installation costs. The firm will straight-line depreciate the equipment to zero using a 5-year recovery period. Additional sales from the overhaul should amount to $238,945.00 per year, and additional operating expenses and other costs (excluding depreciation) will amount to 40.00% of the additional sales. The firm has an ordinary tax rate of 34.00%. What is the operating cash flow for year 1 of this project?
Answer #1

Cost of Equipment = $ 4.15 million or $ 4150000, Installation Cost = $ 222493, Total Initial Cost = 222493 + 4150000 = $ 4372493

Project Tenure = 5 years, Depreciation Format: Straight Line

Annual Depreciation Expense = 4372493 / 5 = $ 874498.6

Additional Sales = $ 238945

Less: Operational Expenses = 40% of Sales = 0.4 x 238495 = $ 95398

EBITDA = $ 143097

Less: Depreciation = $ 874498.6

EBIT = - $ 731401.6

Less: Interest Expenses = $ 0

Profit Before Tax (PBT) = - $ 731401.6

Less: Tax Expenses @ 34 % = $ 0 (as PBT is negative)

Net Income = - $ 731401.6

Add: Annual Depreciation Expense = $ 874498.6

Operating Cash Flow (OCF) = $ 143097